# Quantum dispersionless KdV hierarchy revisited

### v2

@@ 1. Introduction
Since the proof of the Witten-Kontsevich theorem [15, 23], integrable hierarchies have provided important tools for studying various problems in mathematical physics. For example, in 2d topological field theories, corresponding partition functions are controlled by evolutionary PDEs [1, 6]. In symplectic field theories [7], partition functions satisfy a family of Schrödinger type equations, and the commuting quantum Hamiltonian operators can be viewed as quantum integrable hierarchies. While the relationship between 2d topological field theories and evolutionary integrable PDEs have been studied in great detail during last three decades, the interplay between symplectic field theories and quantum integrable hierarchies are still largely unknown.
In [2], a quantization framework is established for arbitrary cohomological field theories, where quantum Hamiltonian structures are given by the canonical quantization of the original Poisson structures of hydrodynamic type, and quantum Hamiltonians are given by intersection numbers and Hodge integrals over double ramification cycles. Let us take the trivial one-dimensional cohomological field theory as an example (see [2, 5, 7]). The genus zero (classical) integrable hierarchy is the dispersionless KdV hierarchy which admits the following Hamiltonian form:
@@
and the corresponding Hamiltonian structure is given by the Poisson bracket of the form
(1.1) $\{v(x),v(y)\}=\delta^{\prime}(x-y).$
Note that the Hamiltonians are in involution:
@@
To quantize this Poisson bracket, we expand the field $v(x)$ into Fourier modes
$v(x)=\sum_{k\in\mathbb{Z}}v_{k}e^{ikx},$
then the Poisson bracket above is equivalent to
$\{v_{m},v_{n}\}=i\,m\,\delta_{m+n,0}.$
By canonical quantization, for $n\geq 1$, let us denote by $\hat{q}_{n}$ the creators corresponding to modes $v_{n}$ and $\hat{p}_{n}$ the annihilators corresponding to modes $v_{-n}$. Then the Poisson bracket is quantized into the commutators
(1.2) $[\hat{p}_{n},\hat{q}_{m}]=\hbar\,n\,\delta_{m,n},\quad[\hat{p}_{n},\hat{p}_{m}]=[\hat{q}_{n},\hat{q}_{m}]=0,\quad n,m\geq 1,$
where $\hbar$ is a formal parameter, and the Hamiltonians $H_{n}^{c}$ is quantized into
$\hat{H}_{n}^{c}=\int:\frac{v^{n+1}}{(n+1)!}:.$
To explicitly compute these quantum Hamiltonians, we first substitute
$v=\sum_{n\geq 1}\hat{q}_{n}e^{inx}+\hat{p}_{n}e^{-inx},$
and normally order these operators by putting annihilators $\hat{p}_{n}$ on the right to obtain
$:\frac{v^{n+1}}{(n+1)!}:\,=\sum_{k\in\mathbb{Z}}f_{k}^{(n)}(\hat{q}_{i},\hat{p}_{j})e^{ikx},$
then we set $\hat{H}_{n}^{c}=f_{0}^{(n)}$. For example, we see that
$\hat{H}_{1}^{c}=\sum_{n\geq 1}\hat{q}_{n}\hat{p}_{n},\quad\hat{H}_{2}^{c}=\frac{1}{2}\sum_{m,n\geq 1}\hat{q}_{n}\hat{q}_{m}\hat{p}_{m+n}+\hat{q}_{m+n}\hat{p}_{n}\hat{p}_{m}.$
One can verify that classically commutative Hamiltonians $H_{n}^{c}$ do not quantize into commutative operators $\hat{H}_{n}^{c}$ under the commutators defined in (1.2). It is observed by Eliashberg that there exist quantum corrections to these operators such that they become mutually commuting operators after corrections. These corrected quantum Hamiltonians are given by the generating series (see [5, 20])
$\hat{H}^{q}(z)=\frac{1}{s(\sqrt{\hbar}z)}\int:\exp\left(zs(i\sqrt{\hbar}z\partial_{x})v(x)\right):,\quad\hat{H}^{q}(z)=1+\sum_{n\geq 0}\hat{H}_{n}^{q}z^{n+1},$
where the function
$s(t)=\frac{e^{t/2}-e^{-t/2}}{t}=\sum_{n\geq 0}\frac{t^{2n}}{4^{n}(2n+1)!}.$
Then it is proved that $\hat{H}_{n}^{q}$ is a correction of $\hat{H}_{n}^{c}$ in the sense that $\hat{H}_{n}^{q}|_{\hbar=0}=\hat{H}_{n}^{c}$ and
$\left[\hat{H}_{n}^{q},\hat{H}_{m}^{q}\right]=0,\quad n,m\geq 1.$
In [2], it is proved that these quantum corrections can be obtained by computing Hodge integrals over double ramification cycles. Furthermore, the corresponding symplectic field theoretic potential is computed in [5]. More aspects of the spectral problem are studied in [10, 11, 21].
In this paper, we approach the quantization problem of the dispersionless KdV hierarchy in the view point of vertex algebra. We obtain a slightly different deformation quantization of the Hamiltonian structure (1.1) from that obtained in [2], and we find a natural combinatorial expression for the corrected quantum Hamiltonians, which indicates a deep relationship between quantum integrable systems and vertex algebra.
To present the main results of the paper, we first introduce some notations. Let us denote by
$\mathcal{A}=\mathbb{C}\left[v,v^{(1)},v^{(2)},\dots\right]$
the ring of differential polynomial and by $\partial_{x}$ the derivation on $\mathcal{A}$ defined by
$\partial_{x}=\sum_{n\geq 0}v^{(n+1)}\frac{\partial}{\partial v^{(n)}},\quad v^{(0)}:=v.$
Using this notation, we can formally view $v$ as a function $v(x)$ with the independent variable $x$, then we have $v^{(n)}=\partial_{x}^{n}v(x)$. Denote by $\mathcal{F}=\mathcal{A}/\partial_{x}\mathcal{A}$ the space of local functionals, and for an element $f\in\mathcal{A}$, denote by $\int f$ its class in $\mathcal{F}$. These spaces play very important roles in the theory of integrable hierarchies. For example, the Hamiltonian structure (1.1) can be written as a Lie bracket on $\mathcal{F}$:
(1.3) $\left\{\int f,\int g\right\}=\int\frac{\delta f}{\delta v}\partial_{x}\frac{\delta g}{\delta v},\quad\frac{\delta f}{\delta v}:=\sum_{n\geq 0}(-\partial_{x})^{n}\frac{\partial f}{\partial v^{(n)}}.$
Denote by $V$ the Heisenberg vertex algebra generated by the field $b(z)$, whose modes form the Heisenberg Lie algebra:
$b(z)=\sum_{n\in\mathbb{Z}}b_{(n)}z^{-n-1},\quad\left[b_{(n)},b_{(m)}\right]=\hbar\,n\,\delta_{m+n,0}.$
Let $|0\rangle\in V$ be the vacuum state, then the state space $V$ is a free $\mathbb{C}[\hbar]$-module spanned by states of the form
@@
Hence, we have a natural $\hbar$-linear isomorphism of vector space
@@
(1.4) $[F,G]=\varphi^{-1}\left(\varphi(G)_{(0)}\varphi(F)\right),\quad F,G\in\mathcal{F}[\hbar],$
defines a deformation quantization of the Hamiltonian structure (1.3). In terms of differential polynomials, this bracket is of the form
@@
here $f,g\in\mathcal{A}$ are arbitrary lifts of $F$ and $G$, respectively.
Next, let us construct the quantization of the Hamiltonians of dispersionless KdV hierarchy. By abusing the notation, we use $b$ to denote the state $b_{(-1)}|0\rangle$. We then define the quantum Hamiltonian densities $h_{n}\in\mathcal{A}[\hbar]$ by the recursion relation
$h_{0}=\varphi^{-1}(b),\quad h_{n}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{1}{\binom{n+1}{k+1}}\varphi^{-1}\left(:\varphi(h_{k})\varphi(h_{n-k-1}):\right),\quad n\geq 1.$
We call this a non-associative Weyl quantization since each $h_{n}$ corresponds to the state obtained by averaging all the possible ways of associating $n+1$ states of $b$. See Sect. 3.1 for details and a combinatorial way for constructing $h_{n}$. Though these Hamiltonian densities are defined via vertex algebra, they satisfy a straightforward recursion relation as differential polynomials.
@@
From this recursion we see that $h_{n}$ is indeed a deformation of the classical Hamiltonians in the sense that
$\lim_{\hbar\to 0}h_{n}=\frac{v^{n+1}}{(n+1)!}.$
@@
$\left[H_{n},H_{m}\right]=0,\quad n,m\geq 0,$
here the bracket is the deformed quantum Hamiltonian structure defined in (1.4). Hence, the quantum Hamiltonians $H_{n}$ form a quantization of the dispersionless KdV hierarchy.
Finally, let us consider the eigenvalue problem. Due to our definition of $h_{n}$, it is natural to consider the operators $\varphi(h_{n})_{(n)}\in\mathrm{End}(V)$.
Remark 1.1.
Note that $\varphi(h_{n})$ is of conformal weight $n+1$, hence the operator $\varphi(h_{n})_{(n)}$ is of conformal weight 0.
@@
Let $f\in\mathcal{A}$ be a differential polynomial. Then for $n\geq 0$, we have
@@
$\cdot\partial_{x}^{2m-1-n+\sum r_{i}+\sum s_{i}}\left(\frac{\partial^{m}f}{\partial v^{(r_{1})}\dots\partial v^{(r_{m})}}\right)\frac{\partial^{m}}{\partial v^{(s_{1})}\dots\partial v^{(s_{m})}}$
[…]
here $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$ is a partition of length $\ell(\lambda)$, and ${\bf{e}}_{n}\in\mathcal{A}[\sqrt{\hbar}]$ is given by
@@
Remark 1.2.
One needs to introduce $\sqrt{\hbar}$ to consider the eigenvalue problem.
@@
Schur polynomials serve as a complete set of common eigenvectors of operators $\hat{H}_{n}$. More precisely, we have
$\hat{H}_{n}S_{\lambda}=\frac{\hbar^{\frac{n+1}{2}}}{n+1}\sum_{k=0}^{n}\binom{n+1}{k}\left(\binom{-\ell(\lambda)}{k+1}+\sum_{i=1}^{\ell(\lambda)}\binom{\lambda_{i}-i}{k}\right)S_{\lambda}.$
As a byproduct, we also prove an explicit formula of $h_{n}$ represented by Schur polynomials.
@@
This paper is organized as follows. In Sect. 2, after recalling some basic facts about vertex algebras, we prove Theorem 1.1 and Proposition 1.4 in a slightly more general setting. In Sect. 3 we construct a quantization of the dispersionless KdV hierarchy and prove various properties of the quantum Hamiltonians. Our main tool for the proofs is based on the work [18] studying actions of $W$-type operators on Schur functions. Finally, in Sect. 4, we give some concluding remarks and compare our results to those obtained in [2, 5].

### v3

@@ 1. Introduction
Since the proof of the Witten-Kontsevich theorem [16, 24], integrable hierarchies have provided important tools for studying various problems in mathematical physics. For example, in 2d topological field theories, the partition functions are controlled by integrable hierarchies of evolutionary Hamiltonian PDEs [2, 7]. In symplectic field theories [8], the partition functions satisfy a family of Schrödinger type equations. The quantum Hamiltonians involved in these Schrödinger equations mutually commute, and therefore can be viewed as quantum integrable hierarchies. While Hamiltonian integrable hierarchies as well as their relationships to 2d topological field theories have been studied in great detail during the last three decades, the corresponding quantum counterparts are still under investigation.
An important advancement toward a general theory of quantum integrable hierarchies was achieved by Buryak and Rossi [3], where a quantization framework is established for arbitrary cohomological field theories using the geometry of double ramification cycles. In their constructions, quantum Hamiltonian structures are given by the canonical quantization of the classical Hamiltonian structures of hydrodynamic types, and quantum Hamiltonians are defined through the intersection theory which is specified by the given cohomological field theory. In particular, their work revealed very deep relationships between quantum integrable hierarchies and geometries of the moduli space of stable curves.
This paper approaches quantum integrable hierarchies in terms of vertex algebras by studying the quantization of the dispersionless KdV hierarchy. The quantization of such a particular integrable hierarchy was addressed both in the framework of symplectic field theory [8, 21] and of the double ramification cycle theory [3], and its properties were studied in a series of works [6, 11, 12, 22]. In the present paper, however, a quantization framework different from above-mentioned ones is established. In particular, we realize a quantum Hamiltonian structure of hydrodynamic type as the Lie algebra canonically attached to the Heisenberg vertex algebra, and apply it to study the quantization of the dispersionless KdV hierarchy. As a result, we construct an Abelian Lie subalgebra consisting of certain states in the Heisenberg vertex algebra, and classical limits of these states coincide with classical Hamiltonian densities of the dispersionless KdV hierarchy. These states provide a quantization of the classical hierarchy, and their properties are studied in detail.
It should be emphasized that the quantum dispersionless KdV hierarchy constructed in this paper looks different from the one studied in the above-mentioned papers. Both the quantum Hamiltonian structure and the quantum Hamiltonians are different. However, despite these differences, the quantization framework based on vertex algebras shares great similarities with the framework based on geometries of moduli spaces. A comparison between these two constructions is given in Sect. 4.
Let us illustrate the basic concepts regarding quantization of Hamiltonian integrable hierarchies and present main results of this paper. We begin by introducing necessary notations. Define the ring of differential polynomial
$\mathcal{A}=\mathbb{C}\left[v,v^{(1)},v^{(2)},\dots\right],$
and denote by
$\partial_{x}=\sum_{n\geq 0}v^{(n+1)}\frac{\partial}{\partial v^{(n)}},\quad v^{(0)}:=v,$
a derivation on $\mathcal{A}$. It follows that we can formally view $v$ as a function $v(x)$ with the independent variable $x$, and we have $v^{(n)}=\partial_{x}^{n}v(x)$. Furthermore, we introduce the quotient space $\mathcal{F}=\mathcal{A}/\partial_{x}\mathcal{A}$ whose elements are called local functionals. For an element $f\in\mathcal{A}$, denote by $F=\int f$ its class in $\mathcal{F}$, and we call $f$ a density of $F$. Using these notions, the (classical) dispersionless KdV hierarchy can be presented by
@@
where the variational derivative is defined by
$\frac{\delta F}{\delta v}:=\sum_{n\geq 0}(-\partial_{x})^{n}\frac{\partial f}{\partial v^{(n)}},\quad F=\int f\in\mathcal{F}.$
Note that the variational derivative of a local functional is independent of the choice of its density.
One of the most important properties of the dispersionless KdV hierarchy is that it is Hamiltonian. We define the classical Hamiltonian structure of the dispersionless KdV hierarchy to be the Lie algebra structure on $\mathcal{F}$ specified by
(1.1) $\left\{F,G\right\}=\int\frac{\delta F}{\delta v}\partial_{x}\frac{\delta G}{\delta v}.$
For a given local functional $F$, this Lie bracket induces a derivation on $\mathcal{A}$ given by
$\{-,F\}\colon\mathcal{A}\to\mathcal{A},\quad\{g,F\}=\sum_{n\geq 0}\frac{\partial g}{\partial v^{(n)}}\partial_{x}^{n+1}\frac{\delta F}{\delta v},\quad g\in\mathcal{A}.$
It follows from the definition that for any $F,G\in\mathcal{F}$ and any density $g\in\mathcal{A}$ of $G$, we have
$\int\{g,F\}=\{G,F\},$
hence we abuse the notation and use the same bracket to denote both the Lie bracket on $\mathcal{F}$ and its induced action on $\mathcal{A}$. Using these notations, the dispersionless KdV hierarchy can be represented by the following Hamiltonian form:
$\frac{\partial v}{\partial t_{n}}=\{v,H_{n}^{c}\}.$
We call $H_{n}^{c}$ the classical Hamiltonians of the dispersionless KdV hierarchy. They are in involution, namely they span an Abelian Lie subalgebra of $\mathcal{F}$:
@@
The Hamiltonian property of the dispersionless KdV hierarchy enables us to formulate its deformation quantization as follows:
  1. (1)
    Construct a Lie algebra bracket [,][-,-] on $\mathcal{F}[\hbar]$ which deforms the classical Hamiltonian structure in the sense that
    $\lim_{\hbar\to 0}\frac{1}{\hbar}[F,G]=\{F,G\},\quad F,G\in\mathcal{F}.$
    This Lie algebra structure is called the quantum Hamiltonian structure.
  2. (2)
    Deform the classical Hamiltonians $H_{n}^{c}$ into quantum Hamiltonians $H_{n}\in\mathcal{F}[\hbar]$ such that
    $H_{n},H_{m}]=0,\quad\lim_{\hbar\to 0}H_{n}=H_{n}^{c},\quad n,m\geq 0.$
    We call the quantum Hamiltonians $H_{n}$ as well as the quantum Hamiltonian structure a quantization of the dispersionless KdV hierarchy.
  3. (3)
    Solve the corresponding Schrödinger equations, i.e., determine a representation of the Lie algebra $\mathcal{F}[\hbar]$ and study the eigenvalue problem of $H_{n}$.
In this paper, we study all these three procedures in a unified framework based on the theory of vertex algebras. Let us denote by $V$ the state space of the Heisenberg vertex algebra generated by the state $b\in V$, whose precise definition and related properties will be given in Sect. 2.1 (or see Chapter 2 of [9]). Associated with $b$, there is a formal power series of operators forming a basis of the Heisenberg Lie algebra:
$Y(b,z)=\sum_{n\in\mathbb{Z}}b_{(n)}z^{-n-1},\quad b_{(n)}\in\mathrm{End}(V),\quad\left[b_{(n)},b_{(m)}\right]=\hbar\,n\,\delta_{m+n,0},$
here the Lie bracket is the commutator on $\mathrm{End}(V)$. Denote by $|0\rangle\in V$ the vacuum state, then the state space $V$ is a free $\mathbb{C}[\hbar]$-module spanned by states of the form
@@
For each state $a\in V$, it follows from the definition of a vertex algebra that we have a corresponding field $Y(a,z)$ given by
$Y(a,z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1},\quad a_{(n)}\in\mathrm{End}(V),$
where $a_{(n)}$ is called the $n$-th mode of $a$. In particular, the $n$-th mode of the state $b$ is exactly $b_{(n)}$. The general formula for $a_{(n)}$ in terms of modes of $b$ is given in Prop. 2.2. We view the unknown function $v(x)$ of the dispersionless KdV hierarchy as a classical field, and we treat $Y(b,z)$ as the corresponding quantum field. It follows that there is a natural $\hbar$-linear isomorphism of vector spaces
@@
Let $T\in\mathrm{End}(V)$ be the infinitesimal translation operator of the Heisenberg vertex algebra, which is characterized by
$T|0\rangle=0,\quad\left[T,b_{(n)}\right]=-nb_{(n-1)},\quad n\in\mathbb{Z}.$
Hence, it follows that
@@
where the action of $\partial_{x}$ is extended to $\mathcal{A}[\hbar]$ by $\partial_{x}\hbar=0$. As a result, the isomorphism $\varphi$ induces a natural isomorphism between the quotient spaces $\mathcal{F}[\hbar]$ and $V/TV$ which we still denote by $\varphi$.
Using the above notations, we present here the main results of the paper. First, we construct a quantum Hamiltonian structure based on the vertex algebra structure.
@@
(1.2) $F,G]=\varphi^{-1}\left(\varphi(G)_{(0)}\varphi(F)\right),\quad F,G\in\mathcal{F}[\hbar],$
defines a deformation quantization of the Hamiltonian structure (1.1), here $\varphi(G)_{(0)}$ is the zeroth mode of the state $\varphi(G)$. In terms of differential polynomials, this bracket is of the form
@@
here $f,g\in\mathcal{A}$ are arbitrary densities of $F$ and $G$, respectively.
Next, we construct the quantum Hamiltonians $H_{n}$ for the quantum dispersionless KdV hierarchy. Let us denote by $h_{n}\in\mathcal{A}[\hbar]$ the differential polynomials determined by the recursion relation
$h_{0}=\varphi^{-1}\left(b_{(-1)}|0\rangle\right),\quad h_{n}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{1}{\binom{n+1}{k+1}}\varphi^{-1}\left(\varphi(h_{k})_{(-1)}\varphi(h_{n-k-1})\right),\quad n\geq 1.$
We call this a non-associative Weyl quantization. Very roughly speaking, $h_{n}$ is obtained by averaging the non-associativity of the Heisenberg vertex algebra. See Sect. 3.1 for details and a combinatorial way for constructing $h_{n}$. Though these quantum Hamiltonians are defined via the Heisenberg vertex algebra, it turns out that they can be computed effectively.
@@
It follows that differential polynomials $h_{n}$ indeed define quantum Hamiltonians, and hence we construct a quantum dispersionless KdV hierarchy.
@@
$\lim_{\hbar\to 0}H_{n}=H_{n}^{c},\quad\left[H_{n},H_{m}\right]=0,\quad n,m\geq 0,$
here the bracket is the quantum Hamiltonian structure defined in (1.2).
Finally, we consider the eigenvalue problem. As discussed before, we need to fix a representation of the quantum Hamiltonian structure. One of the motivations for constructing quantum dispersionless KdV hierarchy using a vertex algebra is the observation that any vertex algebra has a natural representation on its state space. Indeed, for any state $a\in V$, their modes $a_{(n)}$ are already endomorphisms on $V$. Since the state space is identified with the space $\mathcal{A}[\hbar]$ of differential polynomials, it follows that there is a family of actions on $\mathcal{A}[\hbar]$ for any given $f\in\mathcal{A}[\hbar]$, as described by the following proposition.
@@
Let $f\in\mathcal{A}[\hbar]$ be a differential polynomial. Then for $n\geq 0$, the $n$-th mode of the state $\varphi(f)$ induces an endomorphism on $\mathcal{A}[\hbar]$ which reads
@@
$\cdot\partial_{x}^{2m-1-n+\sum r_{i}+\sum s_{i}}\left(\frac{\partial^{m}f}{\partial v^{(r_{1})}\dots\partial v^{(r_{m})}}\right)\frac{\partial^{m}}{\partial v^{(s_{1})}\dots\partial v^{(s_{m})}}.$
It follows that we can associate a family of operators for each $h_{n}$. To consider the eigenvalue problem it is essential to choose operators
$\hat{H}_{n}=\varphi^{-1}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi(h_{n})_{(n)}\mathchoice{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.86108pt}{$\scriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}{\mathrel{\raisebox{0.43057pt}{$\scriptscriptstyle\circ$}}}\varphi\in\mathrm{End}(\mathcal{A})[\hbar]$
corresponding to the $n$-th mode of the state $\varphi(h_{n})$ for $n\geq 0$, and the reason for this specific choice is explained in Sect. 3.3.3. We show that operators $\hat{H}_{n}$ mutually commute with each other, and their common eigenvectors are given by Schur polynomials, which are defined by
$S_{\lambda}=\det{\left({\bf{e}}_{\lambda_{i}-i+j}\right)_{1\leq i,j\leq\ell(\lambda)}}\in\mathcal{A}[\sqrt{\hbar}],$
here $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$ with $\lambda_{1}\geq\dots\geq\lambda_{\ell(\lambda)}>0$ is an integer partition of length $\ell(\lambda)$, and ${\bf{e}}_{n}\in\mathcal{A}[\sqrt{\hbar}]$ is determined by the generating series
@@
Schur polynomials $S_{\lambda}$, where $\lambda$ runs over all integer partitions, serve as a complete set of common eigenvectors of operators $\hat{H}_{n}$. More precisely, we have
$\hat{H}_{n}S_{\lambda}=\frac{\hbar^{\frac{n+1}{2}}}{n+1}\sum_{k=0}^{n}\binom{n+1}{k}\left(\binom{-\ell(\lambda)}{k+1}+\sum_{i=1}^{\ell(\lambda)}\binom{\lambda_{i}-i}{k}\right)S_{\lambda},\quad\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)}).$
In studying these properties, we also arrive at the following explicit formula relating $h_{n}$ to Schur polynomials:
@@
This paper is organized as follows. In Sect. 2, after recalling some basic facts about vertex algebras, we prove Theorem 1.1 and Proposition 1.4 in a slightly more general setting. In Sect. 3 we construct quantum Hamiltonians of the dispersionless KdV hierarchy and prove various properties. Our main tool for the proofs is based on the work [19] studying actions of $W$-type operators on Schur functions. Finally, in Sect. 4, we give some concluding remarks and compare our results to those obtained before [3, 6].